# Demystifying amortized causal discovery with transformers

### v2

@@ Abstract
Supervised learning for causal discovery from observational data often achieves competitive performance despite seemingly avoiding the explicit assumptions that traditional methods require for identifiability. In this work, we analyze CSIvA (Ke et al., 2023b) on bivariate causal models, a transformer architecture for amortized inference promising to train on synthetic data and transfer to real ones. First, we bridge the gap with identifiability theory, showing that the training distribution implicitly defines a prior on the causal model of the test observations: consistent with classical approaches, good performance is achieved when we have a good prior on the test data, and the underlying model is identifiable. Second, we find that CSIvA can not generalize to classes of causal models unseen during training: to overcome this limitation, we show that learning on datasets generated from different types of causal models, unambiguously identifiable in isolation, improves the test generalization. We analyze this empirical evidence with theory, illustrating that the ambiguous cases resulting from the mixture of identifiable causal models are unlikely to occur. Overall, we find that amortized causal discovery still adheres to identifiability theory, violating the previous hypothesis from Lopez-Paz et al. (2015) that supervised learning methods could overcome its restrictions.
@@ 1 Introduction
Causal discovery aims to uncover the underlying causal relationships between variables of a system from pure observations, which is crucial for answering interventional and counterfactual queries when experimentation is impractical or unfeasible (Peters et al., 2017, Pearl, 2009, Spirtes, 2010). Unfortunately, causal discovery is inherently ill-posed (Glymour et al., 2019): unique identification of causal directions requires restrictive assumptions on the class of structural causal models (SCMs) that generated the data (Shimizu et al., 2006, Hoyer et al., 2008, Zhang & Hyvärinen, 2009). These theoretical limitations often render existing methods inapplicable, as the underlying assumptions are usually untestable or difficult to verify in practice (Montagna et al., 2023a).
@@
Recently, supervised learning algorithms trained on synthetic data have been proposed to overcome the need for specific hypotheses, which restrains the application of classical causal discovery methods to real-world problems (Ke et al., 2023b, Lopez-Paz et al., 2015, Li et al., 2020, Lippe et al., 2022, Lorch et al., 2022). Seminal work from Lopez-Paz et al. (2015) argues that this learning-based approach to causal discovery would allow dealing with complex data-generating processes and would greatly reduce the need for explicitly crafting identifiability conditions a-priori: despite this ambitious goal, the output of these methods is generally considered unreliable, as no theoretical guarantee is provided. A pair of non-identifiable structural causal models can be associated with different causal directed acyclic graphs (DAGs) $\mathcal{G}\neq\tilde{\mathcal{G}}$, while entailing the same joint distribution $p$ on the system’s variables. It is thus unclear how a learning algorithm presented with observational data generated from $p$ would be able to overcome these theoretical limits and correctly identify a unique causal structure. However, the available empirical evidence seems not to reflect impossibility results, as these methods yield surprising generalization abilities on several synthetic benchmarks. Our work aims to bridge this gap by studying the performance of a transformer architecture for causal discovery through the lens of the theory of identifiability from observational data. Specifically, we analyze the CSIvA (Causal Structure Induction via Attention) model for causal discovery (Ke et al., 2023b), focusing on bivariate graphs, as they offer a controlled yet non-trivial setting for the investigation. As our starting point, we provide closed-form examples that identify the limitations of CSIvA in recovering causal structures of linear non-Gaussian and nonlinear additive noise models, which are notably identifiable, and demonstrate the expected failures through empirical evidence. These findings suggest that the class of structural causal models that can be identified by CSIvA is inherently dependent on the specific class of SCMs observed during training. Thus, the need for restrictive hypotheses on the data-generating process is intrinsic to causal discovery, both in the traditional and modern learning-based approaches: assumptions on the test distribution either are posited when selecting the algorithm (traditional methods) or in the choice of the training data (learning-based methods). To address this limitation, we theoretically and empirically analyze when training CSIvA on datasets generated by multiple identifiable SCMs with different structural assumptions improves its generalization at test time. Our experimental findings are based on the analysis of $\sim 1$M test runs. In summary:
@@
  • We show that the class of structural causal models that CSIvA can identify is defined by the class of SCMs observed through samples during the training. We reinforce the notion that identifiability in causal discovery inherently requires assumptions, which must be encoded in the training data in the case of learning algorithms for amortized inference: this violates a previous hypothesis in Lopez-Paz et al. (2015), which suggests that these methods could exceed the boundaries of identifiability.
@@ 2 Related works
Out of the scope of this work, there are methods that necessarily require interventional data (Brouillard et al., 2020, Ke et al., 2023a, Scherrer et al., 2022), and learning-based algorithms unsuitable for amortized inference (Lachapelle et al., 2020, Ng et al., 2020, Zheng et al., 2018, Zhang et al., 2022, Bello et al., 2022).
@@ 3 Background and motivation
[…]
@@ 3.1 Causal discovery from observational data
with $f_{2,i}$ invertible map and mutually independent noise terms. When $f_{2,i}$ is a nonlinear function, the post-ANM amounts to the identifiable post-nonlinear model (PNL) (Zhang & Hyvärinen, 2009). When $f_{2,i}$ is the identity function and $f_{1,i}$ nonlinear, it simplifies to the nonlinear additive noise model (ANM)(Hoyer et al., 2008, Peters et al., 2014), which is known to be identifiable, and is described by the set of structural equations:
@@ 3.2 Motivation and problem definition
Causal discovery from observational data relies on specific assumptions, which can be challenging to verify in practice (Montagna et al., 2023a). To address this, recent methods leverage supervised learning for the amortized inference of causal graphs (or simply amortized causal discovery), i.e. optimization of an inference model to directly predict a causal structure from newly provided data (Lopez-Paz et al., 2015, Li et al., 2020, Lippe et al., 2022, Lorch et al., 2022, Ke et al., 2023a, Löwe et al., 2020). While these approaches also aim to reduce reliance on explicit identifiability assumptions, they often lack a clear connection to the existing causal discovery theory, making their outputs generally unreliable. We illustrate this limitation through an example.
@@
Our arguments of Example 1 are valid for all learning methods that approximate the conditional distribution over the space of graphs given the input data (Ke et al., 2023b, Lopez-Paz et al., 2015, Li et al., 2020, Lippe et al., 2022, Lorch et al., 2022), and suggest that these algorithms are at most informative about the equivalence class of the causal graph underlying the observations. However, the available empirical evidence does not seem to highlight these limitations, as in practice these methods can infer the true causal DAG on several synthetic benchmarks. Thus, further investigation is necessary if we want to rely on their output in any meaningful sense. In this work, we analyze these "black-box" approaches through the lens of established theory of causal discovery from observational data (causal inference often lacks experimental data, which we do not consider). We study in detail the CSIvA architecture (Ke et al., 2023b) (see Appendix A), a variation of the transformer neural network (Vaswani et al., 2017) for the supervised learning of algorithms for amortized causal discovery. This model is optimized via maximum likelihood estimation, i.e. finding $\Theta$ that minimizes $-\mathbf{E}_{\mathcal{G},\mathcal{D}}[\ln\hat{p}(\mathcal{G}|\mathcal{D};\Theta)]$, where $\hat{p}(\mathcal{G}|\mathcal{D};\Theta)$ is the conditional distribution of a graph $\mathcal{G}$ given a dataset $\mathcal{D}$ parametrized by $\Theta$. We limit the analysis to CSIvA as it is a simple yet competitive end-to-end approach to learning causal models. While this is clearly a limitation of the paper, our theoretical and empirical conclusions exemplify both the role of theoretical identifiability in modern approaches and the new opportunities they provide. Additionally, it fits well within a line of works arguing that specifically transformers can learn causal concepts Jin et al. (2024), Zhang et al. (2024), Scetbon et al. (2024) and can be explicitly trained to identify different assumptions in context (Gupta et al., 2023).
@@ 4 Experimental results through the lens of theory
In this section, we present a comprehensive analysis of bivariate causal discovery with transformers and its relation to the theoretical boundaries of causal discovery from observational data. We show that suitable assumptions must be encoded in the training distribution to ensure the identifiability of the test data, and we additionally study the effectiveness of training on mixtures of causal models to overcome these limitations, improving generalization abilities. In Appendix C we discuss how our findings can be extended to the case of multivariate causal models.
@@ 4.1 Experimental design
We concentrate our research on causal models of two variables, causally related according to one of the two graphs $X\rightarrow Y$, $Y\rightarrow X$. Bivariate models are the simplest non-trivial setting with a well-known theory of causality inference (Hoyer et al., 2008, Zhang & Hyvärinen, 2009, Peters et al., 2014), but also amenable to manipulation. This allows for comprehensive training and analysis of diverse SCMs and facilitates a clear interpretation of the results.
@@
Unless otherwise specified, in our experiments we train CSIvA on a sample of $15000$ synthetically generated datasets, consisting of $1500$ i.i.d. observations. Each dataset is generated according to a single class of SCMs, defined by the mechanism type and the noise terms distribution. The coefficients of the linear mechanisms are sampled in the range $[-3,-0.5]\cup[0.5,3]$, removing small coefficients to avoid close-to-unfaithful effects (Uhler et al., 2012). Nonlinear mechanisms are parametrized according to a neural network with random weights, a strategy commonly adopted in the literature of causal discovery (see Appendix B.2; alternatively, we provide experiments on data generated simulating nonlinear mechanisms by sampling from a Gaussian process, as described in Appendix D.7). The post-nonlinearity of the PNL model consists of a simple map $z\mapsto z^{3}$. Noise terms are sampled from common distributions and a randomly generated density that we call mlp, previously adopted in Montagna et al. (2023a), defined by a standard Gaussian transformed by a multilayer perceptron (MLP) (Appendix B.2). We name these datasets mechanism-noise to refer to their underlying causal model. For example, data sampled from a nonlinear ANM with Gaussian noise are named nonlinear-gaussian. More details on the synthetic data generation schema are found in Appendix B.2. All data are standardized by their empirical variance to remove opportunities to learn shortcuts (Geirhos et al., 2020, Reisach et al., 2021, Montagna et al., 2023b).
@@ 4.2 Warm up: is CSIvA capable of in and out-of-distribution generalization?
First, we investigate the generalization of CSIvA on datasets sampled from the structural casual model that generates the train distribution, with mechanisms and noise distributions fixed between training and testing. We call this in-distribution generalization. The main goal of these experiments is to validate that the performance of our CSIvA implementation is non-trivial. As a benchmark, we present the accuracy of two state-of-the-art approaches from the literature on causal discovery: we consider the DirectLiNGAM and NoGAM algorithms (Shimizu et al., 2011, Montagna et al., 2023c), respectively designed for the inference on LiNGAM and nonlinear ANM generated data [^1] The causal-learn implementation of the PNL algorithm could not perform better than random on our synthetic post-nonlinear data, and we observed that this was due to the sensitivity of the algorithm to the variance scale. So we report the plot of Figure 1(c) without benchmark comparison. We remark that the point of this experiment is not to make any claims on CSIvA being state-of-the-art but to validate that the performance we obtain in our re-implementation is non-trivial. This is clear for PNL, even without comparison.. The results of Figure 1 show that CSIvA can properly generalize to unseen samples from the training distribution: the majority of the trained models present SHD close to zero and comparable to the relative benchmark algorithm.
@@
In practice, we generally do not know the SCM defining the test distribution, so we are interested in CSIvA’s ability to generalize to data sampled from a class of causal models that is unobserved during training. We call this out-of-distribution generalization (OOD). We study OOD generalization to different noise terms, analyzing the network performance on datasets generated from causal models where the mechanisms are fixed with respect to the training, while the noise distribution varies (e.g., given linear-mlp training samples, testing occurs on linear-uniform data). Orthogonally to these experiments, we empirically validate CSIvA’s OOD generalization over different mechanism types (linear, nonlinear, post-nonlinear), while leaving the noise distribution (mlp) fixed across test and training. In Figure 2(a), we observe that CSIvA cannot generalize across the different mechanisms, as the SHD of a network tested on unseen causal mechanisms approximates that of the random baseline. Further, Figure 2(b) shows that out-of-distribution generalization across noise terms does not work reliably, and it is hard to predict when it might occur. We note that these findings are novel in the literature: OOD experiments in the sense we define can not be found in Ke et al. (2023b) in the first place; Li et al. (2020) empirical results are limited to the OOD generalization from linear Gaussian models to linear models with Exponential, Gumbel and Poisson noise. Lorch et al. (2022) analyses generalization from SCMs with Gaussian noise to Laplace and Cauchy distributions, and fixed mechanisms class. The other works on amortized inference in our related works section 2 generally disregard these experiments.
@@
CSIvA generalizes well to test data generated by the same class of SCMs used for training, in line with the findings in Ke et al. (2023b), which validates our implementation and training procedure. However, it struggles when the test data are out-of-distribution, generated by causal models with different mechanism types and noise distributions that those it was trained on. From a practical perspective, this is a relevant finding, given that existing works on amortized causal discovery lacks systematic experiments on the OOD setting. While training on a wider class of SCMs might overcome this limitation, it requires caution. The identifiability of causal graphs indeed results from the interplay between the data-generating mechanisms and noise distribution. However, as we argue in our Example 1, the class of causal models that a supervised learning algorithm can identify is generally not clear. In what follows, we investigate this point and its implications for CSIvA, showing that the identifiability of the test samples can be ensured by imposing suitable assumptions on the class of SCMs generating the training distribution.

### v3

@@ Abstract
Supervised learning for causal discovery from observational data often achieves competitive performance despite seemingly avoiding the explicit assumptions that traditional methods require for identifiability. In this work, we analyze CSIvA (Ke et al., 2023b) on bivariate causal models, a transformer architecture for amortized inference promising to train on synthetic data and transfer to real ones. First, we bridge the gap with identifiability theory, showing that the training distribution implicitly defines a prior on the causal model of the test observations: consistent with classical approaches, good performance is achieved when we have a good prior on the test data, and the underlying model is identifiable. Second, we find that CSIvA can not generalize to classes of causal models unseen during training: to overcome this limitation, we theoretically and empirically analyze when training CSIvA on datasets generated by multiple identifiable causal models with different structural assumptions improves its generalization at test time. Overall, we find that amortized causal discovery with transformers still adheres to identifiability theory, violating the previous hypothesis from Lopez-Paz et al. (2015) that supervised learning methods could overcome its restrictions.
@@ 1 Introduction
Causal discovery aims to uncover the underlying causal relationships between variables of a system from pure observations, which is crucial for answering interventional and counterfactual queries when experimentation is impractical or unfeasible (Peters et al., 2017; Pearl, 2009; Spirtes, 2010). Unfortunately, causal discovery is inherently ill-posed (Glymour et al., 2019): unique identification of causal directions requires restrictive assumptions on the class of structural causal models (SCMs) that generated the data (Shimizu et al., 2006; Hoyer et al., 2008; Zhang and Hyvärinen, 2009). These theoretical limitations often render existing methods inapplicable, as the underlying assumptions are usually untestable or difficult to verify in practice (Montagna et al., 2023a).
@@
Recently, supervised learning algorithms trained on synthetic data have been proposed to overcome the need for specific hypotheses, which restrains the application of classical causal discovery methods to real-world problems (Ke et al., 2023b; Lopez-Paz et al., 2015; Li et al., 2020; Lippe et al., 2022; Lorch et al., 2022). Seminal work from Lopez-Paz et al. (2015) argues that this learning-based approach to causal discovery would allow dealing with complex data-generating processes and would greatly reduce the need for explicitly crafting identifiability conditions a-priori: despite this ambitious goal, the output of these methods is generally considered unreliable, as no theoretical guarantee is provided. A pair of non-identifiable structural causal models can be associated with different causal directed acyclic graphs (DAGs) $\mathcal{G}\neq\tilde{\mathcal{G}}$, while entailing the same joint distribution $p$ on the system’s variables. It is thus unclear how a learning algorithm presented with observational data generated from $p$ would be able to overcome these theoretical limits and correctly identify a unique causal structure. However, the available empirical evidence seems not to reflect impossibility results, as these methods yield surprising generalization abilities on several synthetic benchmarks. Our work aims to bridge this gap by studying the performance of a transformer architecture for causal discovery through the lens of the theory of identifiability from observational data. Specifically, we analyze the CSIvA (Causal Structure Induction via Attention) model for causal discovery (Ke et al., 2023b), focusing on bivariate graphs, as they offer a controlled yet non-trivial setting for the investigation. As our starting point, we provide closed-form examples that identify the limitations of CSIvA in recovering causal structures of linear non-Gaussian and nonlinear additive noise models, which are notably identifiable, and demonstrate the expected failures through empirical evidence. These findings suggest that the class of structural causal models that can be identified by CSIvA is inherently dependent on the specific class of SCMs observed during training. Thus, the need for restrictive hypotheses on the data-generating process is intrinsic to causal discovery, both in the traditional and modern learning-based approaches: assumptions on the test distribution either are posited when selecting the algorithm (traditional methods) or in the choice of the training data (learning-based methods). To address this limitation, we theoretically and empirically analyze when training CSIvA on datasets generated by multiple identifiable SCMs with different structural assumptions improves its generalization at test time. Our experimental findings are based on the analysis of $\sim\hskip-2.84526pt1$M test runs. In summary:
@@
  • We empirically show that the class of structural causal models that CSIvA can identify is defined by the class of SCMs observed through samples during the training. We reinforce the notion that identifiability in causal discovery inherently requires assumptions, which must be encoded in the training data in the case of learning algorithms for amortized inference: this violates a previous hypothesis in Lopez-Paz et al. (2015), which suggests that these methods could exceed the boundaries of identifiability.
@@ 2 Related works
Out of the scope of this work, there are methods that necessarily require interventional data (Brouillard et al., 2020; Ke et al., 2023a; Scherrer et al., 2022), and learning-based algorithms unsuitable for amortized inference (Lachapelle et al., 2020; Ng et al., 2020; Zheng et al., 2018; Zhang et al., 2022; Bello et al., 2022).
@@ 3 Background and motivation
$X_{i}\coloneqq f_{i}(X_{\operatorname{PA}^{\mathcal{G}}_{i}},N_{i}),\hskip 5.69054pt\forall i=1,\ldots,k.$ (1)
@@ 3.1 Causal discovery from observational data
with $f_{2,i}$ invertible map and mutually independent noise terms. When $f_{2,i}$ is a nonlinear function, the post-ANM amounts to the identifiable post-nonlinear model (PNL) (Zhang and Hyvärinen, 2009). When $f_{2,i}$ is the identity function and $f_{1,i}$ nonlinear, it simplifies to the nonlinear additive noise model (ANM)(Hoyer et al., 2008; Peters et al., 2014), which is known to be identifiable, and is described by the set of structural equations:
@@
The special case of linear Gaussian models, with noise terms with equal variance, is also known to be identifiable (Peters and Bühlmann, 2013).
@@ 3.2 Motivation and problem definition
Causal discovery from observational data relies on specific assumptions, which can be challenging to verify in practice (Montagna et al., 2023a). To address this, recent methods leverage supervised learning for the amortized inference of causal graphs (or simply amortized causal discovery), i.e. optimization of an inference model to directly predict a causal structure from newly provided data (Lopez-Paz et al., 2015; Li et al., 2020; Lippe et al., 2022; Lorch et al., 2022; Ke et al., 2023a; Löwe et al., 2020). While these approaches also aim to reduce reliance on explicit identifiability assumptions, they often lack a clear connection to the existing causal discovery theory, making their outputs generally unreliable. We illustrate this limitation through an example.
@@
The claims made in our Example 1 are valid for all learning methods that approximate the conditional distribution over the space of graphs given the input data (Ke et al., 2023b; Lopez-Paz et al., 2015; Li et al., 2020; Lippe et al., 2022; Lorch et al., 2022), and suggest that these algorithms are at most informative about the equivalence class of the causal graph underlying the observations. However, the available empirical evidence does not seem to highlight these limitations, as in practice these methods can infer the true causal DAG on several synthetic benchmarks. Thus, further investigation is necessary if we want to rely on their output in any meaningful sense. In this work, we analyze these "black-box" approaches through the lens of established theory of causal discovery from observational data (causal inference often lacks experimental data, which we do not consider). We study in detail the CSIvA architecture (Ke et al., 2023b) (described in Appendix A), a variation of the transformer neural network (Vaswani et al., 2017) for the supervised learning of algorithms for amortized causal discovery. This model is optimized via maximum likelihood estimation, i.e. finding $\Theta$ that minimizes $\mathbf{E}_{\mathcal{G},\mathcal{D}}[\ln\hat{p}(\mathcal{G}|\mathcal{D};\Theta)]$, where $\hat{p}(\mathcal{G}|\mathcal{D};\Theta)$ is the conditional distribution of a graph $\mathcal{G}$ given a dataset $\mathcal{D}$ parametrized by $\Theta$. We limit the analysis to CSIvA as it is a simple yet competitive end-to-end approach to learning causal models. While this is clearly a limitation of the paper, our theoretical and empirical conclusions exemplify both the role of theoretical identifiability in modern approaches and the new opportunities they provide. Additionally, it fits well within a line of works arguing that specifically transformers can learn causal concepts Jin et al. (2024); Zhang et al. (2024); Scetbon et al. (2024) and can be explicitly trained to identify different assumptions, such as linearity of the mechanisms or normality of the noise terms, when doing in-context learning (Gupta et al., 2023; Brown et al., 2020).
@@ 4 Experimental results through the lens of theory
In this section, we present a comprehensive analysis of bivariate causal discovery with transformers and its relation to the theoretical boundaries of causal discovery from observational data  [^1] The code for CSIvA implementation can be found here. The code for reproducing the experiments of the paper can be found here. We show that suitable assumptions must be encoded in the training distribution to ensure the identifiability of the test data, and we additionally study the effectiveness of training on mixtures of causal models to overcome these limitations, improving generalization abilities. In Appendix C we discuss how our findings can be naturally extended to the case of multivariate causal models: the intuition is that, in this case, identifiability can be guaranteed by iteratively verifying that the causal order of all bivariate subgraphs is individually identifiable (Theorem 28 in Peters et al. (2014)). As a result, it is common to limit the analyses to bivariate graphs (see e.g. Hoyer et al. (2008); Zhang and Hyvärinen (2009); Immer et al. (2023); Xi et al. (2025)), which justifies our choice.
@@ 4.1 Experimental design
We concentrate our research on causal models of two variables, causally related according to one of the two graphs $X\rightarrow Y$, $Y\rightarrow X$. Bivariate models are the simplest non-trivial setting with a well-known theory of causality inference (Hoyer et al., 2008; Zhang and Hyvärinen, 2009; Peters et al., 2014), but also amenable to manipulation. This allows for comprehensive training and analysis of diverse SCMs and facilitates a clear interpretation of the results.
@@
Unless otherwise specified, in our experiments we train CSIvA on a sample of $15000$ synthetically generated datasets, consisting of $1500$ i.i.d. observations. Classes of SCMs are defined by the mechanism type and the noise terms distribution (e.g., linear non-Gaussian): each dataset is generated from a single SCM instance sampled from that class. The coefficients of the linear mechanisms are sampled in the range $3,-0.5]\cup[0.5,3]$, removing small coefficients to avoid close-to-unfaithful effects (Uhler et al., 2012). Nonlinear mechanisms are parametrized according to a neural network with random weights, a strategy commonly adopted in the literature of causal discovery (see Appendix B.2; alternatively, we provide experiments on data generated simulating nonlinear mechanisms by sampling from a Gaussian process, as described in Appendix D.7). The post-nonlinearity of the PNL model consists of a simple map $z\mapsto z^{3}$. Noise terms are sampled from common distributions and a randomly generated density that we call mlp, previously adopted in Montagna et al. (2023a), defined by a standard Gaussian transformed by a multilayer perceptron (MLP) (Appendix B.2). We name these datasets mechanism-noise to refer to their underlying causal model. For example, data sampled from a nonlinear ANM with Gaussian noise are named nonlinear-gaussian. More details on the synthetic data generation schema are found in Appendix B.2. All data are standardized by their empirical variance to remove opportunities to learn shortcuts (Geirhos et al., 2020; Reisach et al., 2021; Montagna et al., 2023b).
@@ 4.2 Warm up: is CSIvA capable of in and out-of-distribution generalization?
First, we investigate the generalization of CSIvA on datasets sampled from the structural casual model that generates the train distribution, with mechanisms and noise distributions fixed between training and testing. We call this in-distribution generalization. The main goal of these experiments is to validate that the performance of our CSIvA implementation is non-trivial. As a benchmark, we present the accuracy of two state-of-the-art approaches from the literature on causal discovery: we consider the DirectLiNGAM and NoGAM algorithms (Shimizu et al., 2011; Montagna et al., 2023c), respectively designed for the inference on LiNGAM and nonlinear ANM generated data [^2] The causal-learn implementation of the PNL algorithm could not perform better than random on our synthetic post-nonlinear data, and we observed that this was due to the sensitivity of the algorithm to the variance scale. So we report the plot of Figure 1(c) without benchmark comparison. We remark that the point of this experiment is not to make any claims on CSIvA being state-of-the-art but to validate that the performance we obtain in our re-implementation is non-trivial. This is clear for PNL, even without comparison.. The results of Figure 1 show that CSIvA can properly generalize to unseen samples from the training distribution: the majority of the trained models present SHD close to zero and comparable to the relative benchmark algorithm.
@@
In practice, we generally do not know the SCM defining the test distribution, so we are interested in CSIvA’s ability to generalize to data sampled from a class of causal models that is unobserved during training. We call this out-of-distribution generalization (OOD). We study OOD generalization to different noise terms, analyzing the network performance on datasets generated from causal models where the mechanisms are fixed with respect to the training, while the noise distribution varies (e.g., given linear-mlp training samples, testing occurs on linear-uniform data). Orthogonally to these experiments, we empirically validate CSIvA’s OOD generalization over different mechanism types (linear, nonlinear, post-nonlinear), while leaving the noise distribution (mlp) fixed across test and training. In Figure 2(a), we observe that CSIvA cannot generalize across the different mechanisms, as the SHD of a network tested on unseen causal mechanisms approximates that of the random baseline. Further, Figure 2(b) shows that out-of-distribution generalization across noise terms does not work reliably, and it is hard to predict when it might occur. We note that these findings are novel in the literature: OOD experiments in the sense we define can not be found in Ke et al. (2023b); Li et al. (2020) empirical results are limited to the OOD generalization from linear Gaussian models to linear models with Exponential, Gumbel and Poisson noise. Lorch et al. (2022) analyses generalization from SCMs with Gaussian noise to Laplace and Cauchy distributions, and fixed mechanisms class. The remaining literature on amortized inference that we discuss in Section˜2 generally disregards these experiments.
@@
CSIvA generalizes well to test data generated by the same class of SCMs used for training, in line with the findings in Ke et al. (2023b), which validates our implementation and training procedure. However, it struggles when the test data are out-of-distribution, generated by causal models with different mechanism types and noise distributions than those it was trained on. From a practical perspective, this is a relevant finding, given that existing works on amortized causal discovery lack systematic experiments on the OOD setting. While training on a wider class of SCMs might overcome this limitation, it requires caution. The identifiability of causal graphs indeed results from the interplay between the data-generating mechanisms and noise distribution. However, as we argue in our Example 1, the class of causal models that a supervised learning algorithm can identify is generally not clear. In what follows, we investigate this point and its implications for CSIvA, showing that the identifiability of the test samples can be ensured by imposing suitable assumptions on the class of SCMs generating the training distribution.